OpenStudy (anonymous):

need integration help: Integral (1/x^2)dx Integral (e^-x)dx Integral (sin(t))(cos(t))dt Integral (radical s) ds

OpenStudy (anonymous):

OK; int(1/x^2)=(-1/x) int(e^-x)=-e^(-x) int (sintcost)=(1/4)sin2t int (radical s)=1/(x^(3/2)) do you need more explanation?

OpenStudy (anonymous):

First integral: reverse power rule Second Integral: U-sub 3rd: U-sub Not sure about the 4th one.

OpenStudy (anonymous):

rewrite the integrands into known forms \[\int{1\over x^2}dx=\int x^{-2}dx\longrightarrow{\rm solve}\\ \int e^{-x}dx\qquad\text{substiture u=-x}\\ \int \sin(t)\cos(t)dt\qquad\text{set }u=\cos(t)\\ \int \sqrt{s}ds=\int s^{1/2}ds\longrightarrow{\rm solve} \]

OpenStudy (anonymous):

\[ \Huge \color{Lime}{~\:\:\:\:\:\:\:\:\:\:\:\:\frak Welcome\:\:\frak To\:\:\frak Open\frak Study} \]